OpenStudy (anonymous):

an=3n-2 start with n=1 and first 5 terms

OpenStudy (anonymous):

a1 =3(1)-2 a1=1 a2= 3(2)-2 a2=4 a2-a1=4-1 d=3 a5=a1+(n-1)d a5=1+(5-1)3 a5=13